OpenStudy (anonymous):

what is the derivative of 4cos2pi(s) and 3sin2pi(s)?

OpenStudy (anonymous):

the derivative of cosine is minus sine so the first one, by the chain rule, is \[-8\pi\sin(2\pi s)\]

OpenStudy (anonymous):

second one is similar, but the derivative of sine is cosine, so again by the chain rule you get \[6\pi\cos(2\pi x)\]